OpenStudy (anonymous):

What is the reciprocal of 2/9 ?

Nnesha (nnesha):

the reciprocal of a/b is b/a just flip the fraction

OpenStudy (astrophysics):

It's actually \[\frac{ 1 }{ \left( \frac{ a }{ b } \right) } = \frac{ b }{ a }\]

OpenStudy (astrophysics):

So all a reciprocal is, \[\frac{ 1 }{ number }\]

jimthompson5910 (jim_thompson5910):

two things to add on 1) if you have a mixed number, convert it to an improper fraction and then use the rule that @Nnesha wrote out 2) The value of 'a' cannot be zero. The value of 'b' cannot be zero. This is to avoid division by zero errors